OpenStudy (anonymous):

the center of the circle defined by 2x^(2)+2y^(2)+4x-12y-30=0 is a) (-1,3) b) (2,-6) c) (-2,6) d) (1,-3)

OpenStudy (anonymous):

I'm guessing I divide everything by 2 to begin with, but what happens to the squares?

OpenStudy (anonymous):

-1,3

OpenStudy (anonymous):

You just have to turn everything into a more usable form : (x-a)^2-(y-b)^2=radius^2, where a and b are the coordinates of the circle's center :) Hope that helped

OpenStudy (anonymous):

adams right

OpenStudy (anonymous):

(x-a)^2+(y-b)^2=r^2 is the form.

OpenStudy (anonymous):

So mainly you should try to complete the squares.

OpenStudy (r):

Do you know how to "complete the squares" for this problem?